how do solve this problem
1-2/3y=9/5-y/5+3/5

Answer by checkley71(8403)

1-2/3Y=9/5-Y/5+3/5
-2/3Y+Y/5=9/5+3/5-1
-2Y/3+Y/5=12/5-1
(-10Y+3Y)/15=7/5
-7Y/15=7/5
-35Y=105
Y=105/-35
Y=-3
PROOF
1-(2/3)-3=9/5-(-3/5)+3/5
1+6/3=9/5+3/5+3/5
1+2=15/5
3=3